# Joint Discrete Precoding and RIS Optimization for RIS-Assisted MU-MIMO Communication Systems

### v1

@@ Abstract
This paper considers a multi-user multiple-input multiple-output (MU-MIMO) system where the downlink communication between a base station (BS) and multiple user equipments (UEs) is aided by a reconfigurable intelligent surface (RIS). We study the sum-rate maximization problem with the objective of finding the optimal precoding vectors and RIS configuration. Due to fronthaul limitation, each entry of the precoding vectors must be picked from a finite set of quantization labels. Furthermore, two scenarios for the RIS are investigated, one with continuous infinite-resolution reflection coefficients and another with discrete finite-resolution reflection coefficients. A novel framework is developed which, in contrast to the common literature that only offers sub-optimal solutions for optimization of discrete variables, is able to find the optimal solution to problems involving discrete constraints. Based on the classical weighted minimum mean square error (WMMSE), we transform the original problem into an equivalent weighted sum mean square error (MSE) minimization problem and solve it iteratively. We compute the optimal precoding vectors via an efficient algorithm inspired by sphere decoding (SD). For optimizing the discrete RIS configuration, two solutions based on the SD algorithm are developed: An optimal SD-based algorithm and a low-complexity heuristic method that can efficiently obtain RIS configuration without much loss in optimality. The effectiveness of the presented algorithms is corroborated via numerical simulations where it is shown that the proposed designs are remarkably superior to the commonly used benchmarks.
@@ I Introduction
A typical 5G base station (BS) is composed of an advanced antenna system (AAS) and a baseband unit (BBU) [7]. AAS is a component consisting of antennas and their corresponding radio units (RUs) and BBU is responsible for processing received uplink and transmitted downlink data. A digital fronthaul with limited capacity connects the AAS and BBU. Due to the capacity limitation of the fronthaul link, the precoding vectors computed at the BBU must be quantized before being sent to the AAS. This makes quantization errors inevitable. Moreover, the literature has reported the design of RIS with either continuous or discrete phase shifts where the former structure is mainly fabricated using varactor diodes and the latter is designed using positive intrinsic negative (PIN) diodes. Specifically, the design of reflective surfaces with continuous phase shifts can be found in [8] where the RIS reflection coefficients can be continuously controlled by changing the voltage applied on the diode. Discrete reflection coefficient implementations with 1-, 2-, and 3-bit resolutions have been presented in [9], [10], and [11], respectively, where changing the reflection coefficient is achieved by turning on/off the PIN diodes. Both continuous and discrete RIS reflection coeffients are practically implementable. However, due to hardware complexity of the continuous phase shift design, using discrete phase shifts at the RIS is more practically appealing, especially when an RIS with a very large number of elements is to be built. Hence, RIS is expected to be of finite resolution in practice, where each of its elements can only select phase shifts from a discrete set and the reflection coefficient of each element accordingly belongs to a discrete set as well.
In this paper, we study a MU-MIMO downlink communication assisted by an RIS, where the direct links between single-antenna UEs and the multi-antenna BS are obstructed. The downlink data is calculated at the BBU and is sent to the AAS through the limited-capacity fronthaul infrastructure. Our objective is to design joint optimal precoding matrix and RIS configuration for sum-rate maximization. This paper aims to find practical downlink precoding and RIS configurations that adhere to the aforementioned hardware constraints.
@@ I-A Related Work
Additionally, there have been some prior works on sum rate maximization in RIS-assisted MU-MIMO systems. Reference [5] investigated the weighted sum rate maximization problem by designing the BS precoding and RIS configuration. They first used the closed-form fractional programming approach to obtain a more tractable problem and then utilized the block coordinate descent algorithm to solve the resultant problem. The weighted sum rate maximization problem in a multi-cell MIMO communication system was studied in [6], where the authors have utilized the weighted minimum mean square error (WMMSE) approach for alternatively optimizing the BS precoding and RIS configuration. In [31], an aerial ground communication scenario was considered where an RIS is deployed to assist the communication between a UAV and multiple ground users. The sum rate maximization problem was studied under this setup aiming to optimize the UAV trajectory and RIS configuration. The authors have employed a conjugate gradient algorithm to optimize the RIS configuration and utilized a successive convex approximation approach to find the UAV trajectory. The authors in [32] studied a sum-rate maximization problem in a multi-user RIS-assisted system by leveraging the statistical CSI at the BS and used the random matrix theory to derive the asymptotic sum rate. They then provided a new formulation of the sum-rate maximization problem with the asymptotic sum-rate being the objective function. After decoupling the precoding and RIS optimization problems, a projected gradient ascent algorithm and a water-filling approach were applied to find sub-optimal RIS configuration and precoding vectors, respectively. All the aforementioned works on sum-rate maximization in RIS-assisted systems only considered continuous RIS phase shifts. In [22], Di et al. considered discrete phase shifts at the RIS and formulated a mixed-integer sum rate maximization problem. They then proposed an iterative approach for designing the BS precoding and RIS configuration. However, as mentioned earlier, this work only provides a sub-optimal solution for RIS configuration. Besides, none of the above works take the fronthaul limitation at the BS side into account.
@@ I-B Contributions
We study the joint precoding and RIS configuration design in practical RIS-assisted MU-MIMO communication systems. We start by assuming a continuous infinite-resolution RIS and then extend our designs to the practical discrete RIS. We utilize an SD-based algorithm to design the optimal precoding vectors and discrete RIS configuration. This proposed framework for optimizing discrete precoding and RIS configurations is a major step toward operating the technology efficiently under practical hardware limitations. The main contributions of this paper our as outlined below:
  • We formulate a sum-rate maximization problem for optimizing the BS precoding vectors and RIS configuration. Two RIS configuration scenarios are considered, one with continuous infinite-resolution RIS reflection coefficients and another with discrete finite-resolution reflection coefficients, where in the latter scenario, only a few reflection coefficients are available for RIS elements to choose from. We take into account the fronthaul limitation for solving the sum rate maximization problems. Though joint precoding and RIS configuration design has been studied in some prior works (e.g., [33]), this is the first work that considers both discrete precoding and finite-resolution RIS configurations. We transform the original sum rate maximization problem into an equivalent weighted sum MSE minimization problem and adopt an alternating optimization approach based on the block coordinate descent method to solve it.
  • We propose a solution based on the Schnorr Euchner SD (SESD) algorithm to solve the precoding optimization sub-problem. This solution provides the global optimal solution to the precoding optimization sub-problem while significantly reducing the complexity compared to an exhaustive search. We further employ the SESD algorithm to solve the RIS configuration optimization problem when RIS reflection coefficients belong to a discrete set. In contrast to the existing literature, where only suboptimal solutions for discrete RIS configuration are offered, the SESD algorithm finds the optimal discrete RIS configurations. This is of paramount importance in a MU-MIMO RIS-assisted system, where suboptimal configuration of the RIS limits the interference suppression capability of the system and severely degrades the performance.
@@
  • We evaluate all the proposed algorithms via numerical simulations under different system setups and show the superiority of the presented SESD-based designs over the commonly used benchmarks. We also assess the effectiveness of the proposed low-complexity heuristic SESD algorithm for RIS configuration optimization and show that it performs close to the SESD design.
@@ III Iterative Weighted Sum MSE Minimization
in the sense that the optimal precoding vectors $\{\mathbf{w}_{k}\}$ and optimal RIS configuration vector $\boldsymbol{\theta}$ are the same for both problems. In problem (9), $\mathbf{c}=[c_{1},\ldots,c_{K}]^{\mbox{\tiny$\mathrm{T}$}}$ and $\boldsymbol{\beta}=[\beta_{1},\ldots,\beta_{K}]^{\mbox{\tiny$\mathrm{T}$}}$.
@@ III-A Optimizing Precoding Vectors
$\mathbf{w}_{k}^{\mathrm{inf}}=\mathbf{V}^{-1}\mathbf{v}_{k}^{*}$ (18)
@@
25:            $\tilde{\mathbf{w}}_{k}[m]\leftarrow\check{\mathbf{w}}_{k}[m]$
@@
$\xi_{m}=\tilde{\mathbf{d}}_{w,k}[m]-\tilde{\mathbf{R}}_{w}[m,m+1:2M]\check{\mathbf{w}}_{k}[m+1:2M]$ (24)
@@
is chosen as the new search radius. Then, fixing $\check{\mathbf{w}}_{k}[3],\ldots,\check{\mathbf{w}}_{k}[2M]$, the algorithm updates $\check{\mathbf{w}}_{k}[2]$ by selecting another label, i.e., $\mathcal{L}(\mathbf{p}_{2}[2])$ and checks if $\mathbf{r}[2]=\sum_{i=2}^{2M}\left|\tilde{\mathbf{d}}_{w,k}[i]-\tilde{\mathbf{R}}_{k}[i,i:2M]\tilde{\mathbf{w}}_{k}[i:2M]\right|^{2}$ with the newly updated $\check{\mathbf{w}}_{k}[2]$ exceeds the search radius $r_{\mathrm{opt}}$ or not. The reason why the algorithm does not update $\check{\mathbf{w}}_{k}[1]$ and directly goes for updating $\check{\mathbf{w}}_{k}[2]$ is that for the combination of entries $\check{\mathbf{w}}_{k}[2],\ldots,\check{\mathbf{w}}_{k}[2M]$, the value $\check{\mathbf{w}}_{k}[1]$ obtained by solving (23) for $m=1$ is the optimal choice among all the quantization labels and results in the smallest possible radius; thus, there is no need to explore other labels in the set at this level. If $\mathbf{r}[2]<r_{\mathrm{opt}}$, the algorithm proceeds to update $\check{\mathbf{w}}_{k}[1]$ by solving (23) again. Otherwise, it moves to an upper level and updates $\check{\mathbf{w}}_{k}[3]$ having $\check{\mathbf{w}}_{k}[4],\ldots,\check{\mathbf{w}}_{k}[2M]$ fixed. This procedure is repeated and after each successful update of $\check{\mathbf{w}}_{k}[m]$ (i.e., the update that results in $\mathbf{r}[m]<r_{\mathrm{opt}}$), the algorithm moves to a lower level and finds $\check{\mathbf{w}}_{k}[m-1]$ from (23). When updating $\check{\mathbf{w}}_{k}[m]$ is unsuccessful, the value of $m$ is increased by 1 and $\check{\mathbf{w}}_{k}[m+1]$ is updated by choosing another label from $\mathcal{L}$ based on the sorting indices at this level. Whenever a new combination of entries $\check{\mathbf{w}}_{k}[1],\ldots,\check{\mathbf{w}}_{k}[2M]$ is found that satisfies $\mathbf{r}[1]=\|\tilde{\mathbf{d}}_{w,k}-\tilde{\mathbf{R}}_{w}\check{\mathbf{w}}_{k}\|^{2}<r_{\mathrm{opt}}$, the search radius is updated as $r_{\mathrm{opt}}=\mathbf{r}[1]$. It is important to note the role of sorting indices in updating the precoding entries. Specifically, assuming that $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i])$, the next time the algorithm moves from level $m-1$ to level $m$, $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$ is selected as the new precoding entry at this level. This is because if $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$ fails to fall in the optimal radius $r_{\mathrm{opt}}$, there is no need for further update of $\check{\mathbf{w}}_{k}[m]$ for the given values of $\check{\mathbf{w}}_{k}[m+1],\ldots,\check{\mathbf{w}}_{k}[2M]$. In other words, if $\mathbf{r}[m]>r_{\mathrm{opt}}$ for $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$, we will surely have $\mathbf{r}[m]>r_{\mathrm{opt}}$ for $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i^{\prime}]),\,i^{\prime}>i+1$.
@@ III-B Optimizing the RIS Configuration
$\bar{\theta}_{\hat{n}}=e^{j\left(-\arg\left(\sum_{n\neq\hat{n}}\theta_{n}^{*}[\mathbf{A}]_{n,\hat{n}}-[\mathbf{t}]^{*}_{\hat{n}}\right)+\pi\right)}$ (34)
@@ III-C Overall Algorithm
7:Given $\overline{\mathbf{W}}^{(0)}$, $\overline{\boldsymbol{\theta}}^{(0)}$, and $\overline{\boldsymbol{\beta}}^{0}$, find $e_{k}^{(0)}$ from (III) and compute $f^{(0)}=\sum_{k=1}^{K}\overline{c}_{k}^{(0)}e_{k}^{(0)}-\ln(\overline{c}_{k}^{(0)})$
@@
10:   Find the optimal precoding vectors $\{\overline{\mathbf{w}}_{k}^{(l)}\}$ from Algorithm 1
@@
15:   Compute $f^{(l)}=\sum_{k=1}^{K}\overline{c}_{k}^{(l)}e_{k}^{(l)}-\ln(\overline{c_{k}}^{(l)})$
@@
IV RIS with Finite Phase-Shift Resolution
In practice, due to hardware design limitations, it may not be possible for RIS elements to take any phase value. Therefore, a practical RIS is expected to have a finite phase-shift resolution, meaning that only a limited number of phase shifts are available for RIS elements to choose from. In such a case, the discrete set of reflection coefficients is given by

### v2

@@ Abstract
This paper considers a multi-user multiple-input multiple-output (MU-MIMO) system where the downlink communication between a base station (BS) and multiple user equipments (UEs) is aided by a reconfigurable intelligent surface (RIS). We study the sum rate maximization problem with the objective of finding the optimal precoding vectors and RIS configuration. Due to fronthaul limitation, each entry of the precoding vectors must be picked from a finite set of quantization labels. Furthermore, two scenarios for the RIS are investigated, one with continuous infinite-resolution reflection coefficients and another with discrete finite-resolution reflection coefficients. A novel framework is developed which, in contrast to the common literature that only offers sub-optimal solutions for optimization of discrete variables, is able to find the optimal solution to problems involving discrete constraints. Based on the classical weighted minimum mean square error (WMMSE), we transform the original problem into an equivalent weighted sum mean square error (MSE) minimization problem and solve it iteratively. We compute the optimal precoding vectors via an efficient algorithm inspired by sphere decoding (SD). For optimizing the discrete RIS configuration, two solutions based on the SD algorithm are developed: An optimal SD-based algorithm and a low-complexity heuristic method that can efficiently obtain RIS configuration without much loss in optimality. The effectiveness of the presented algorithms is corroborated via numerical simulations where it is shown that the proposed designs are remarkably superior to the commonly used benchmarks.
@@ I Introduction
A typical 5G base station (BS) is composed of an advanced antenna system (AAS) and a baseband unit (BBU) [7]. AAS is a component consisting of antennas and their corresponding radio units (RUs) and BBU is responsible for processing received uplink and transmitted downlink data. A digital fronthaul connects the AAS and BBU. The fronthaul connection consists of fiber cables with a limited capacity and needs to transfer a number of digital samples per second that grows with the bandwidth, antennas, and users. Due to the capacity limitation of the fronthaul, each sample must be compressed to be represented by a limited number of bits. Hence, the entries of the precoding vectors computed at the BBU cannot take any value and must belong to a specified discrete set determined by the fronthaul interface. This makes quantization errors inevitable. Moreover, the literature has reported the design of RIS with either continuous or discrete phase shifts where the former structure is mainly fabricated using varactor diodes and the latter is designed using positive intrinsic negative (PIN) diodes. Specifically, the design of reflective surfaces with continuous phase shifts can be found in [8] where the RIS reflection coefficients can be continuously controlled by changing the voltage applied on the diode. Discrete reflection coefficient implementations with 1-, 2-, and 3-bit resolutions have been presented in [9], [10], and [11], respectively, where changing the reflection coefficient is achieved by turning on/off the PIN diodes. Both continuous and discrete RIS reflection coeffients are practically implementable. However, due to hardware complexity of the continuous phase shift design, using discrete phase shifts at the RIS is more practically appealing, especially when an RIS with a very large number of elements is to be built. Hence, RIS is expected to be of finite resolution in practice, where each of its elements can only select phase shifts from a discrete set and the reflection coefficient of each element accordingly belongs to a discrete set as well.
In this paper, we study a MU-MIMO downlink communication assisted by an RIS, where the direct links between single-antenna UEs and the multi-antenna BS are obstructed. The downlink data is calculated at the BBU and is sent to the AAS through the limited-capacity fronthaul infrastructure. Our objective is to design joint optimal precoding matrix and RIS configuration for sum rate maximization. This paper aims to find practical downlink precoding and RIS configurations that adhere to the aforementioned hardware constraints.
@@ I-A Related Work
Additionally, there have been some prior works on sum rate maximization in RIS-assisted MU-MIMO systems. Reference [5] investigated the weighted sum rate maximization problem by designing the BS precoding and RIS configuration. They first used the closed-form fractional programming approach to obtain a more tractable problem and then utilized the block coordinate descent algorithm to solve the resultant problem. The weighted sum rate maximization problem in a multi-cell MIMO communication system was studied in [6], where the authors have utilized the weighted minimum mean square error (WMMSE) approach for alternatively optimizing the BS precoding and RIS configuration. In [31], an aerial ground communication scenario was considered where an RIS is deployed to assist the communication between a UAV and multiple ground users. The sum rate maximization problem was studied under this setup aiming to optimize the UAV trajectory and RIS configuration. The authors have employed a conjugate gradient algorithm to optimize the RIS configuration and utilized a successive convex approximation approach to find the UAV trajectory. The authors in [32] studied a sum rate maximization problem in a multi-user RIS-assisted system by leveraging the statistical CSI at the BS and used the random matrix theory to derive the asymptotic sum rate. They then provided a new formulation of the sum rate maximization problem with the asymptotic sum rate being the objective function. After decoupling the precoding and RIS optimization problems, a projected gradient ascent algorithm and a water-filling approach were applied to find sub-optimal RIS configuration and precoding vectors, respectively. All the aforementioned works on sum rate maximization in RIS-assisted systems only considered continuous RIS phase shifts. In [22], Di et al. considered discrete phase shifts at the RIS and formulated a mixed-integer sum rate maximization problem. They then proposed an iterative approach for designing the BS precoding and RIS configuration. However, as mentioned earlier, this work only provides a sub-optimal solution for RIS configuration. Besides, none of the above works take the fronthaul limitation at the BS side into account.
@@ I-B Contributions
We study the joint precoding and RIS configuration design in practical RIS-assisted MU-MIMO communication systems. We start by assuming a continuous infinite-resolution RIS and then extend our design to the practical discrete RIS. We utilize an SD-based algorithm to design the optimal precoding vectors and discrete RIS configuration. This proposed framework for optimizing discrete precoding and RIS configurations is a major step toward operating the technology efficiently under practical hardware limitations. The main contributions of this paper are as outlined below:
  • We formulate a sum rate maximization problem for optimizing the BS precoding vectors and RIS configuration. Two RIS configuration scenarios are considered, one with continuous infinite-resolution reflection coefficients and another with discrete finite-resolution reflection coefficients, where in the latter scenario, only a few reflection coefficients are available for RIS elements to choose from. We also take into account the fronthaul capacity limitation when solving the sum rate maximization problems, necessitating the precoding entries to be selected from a discrete set. The discrete constraints on the RIS configuration and precoding make the problem a non-convex combinatorial optimization problem. Though joint precoding and RIS configuration design has been studied in some prior works (e.g., [33]), this is the first work that considers both discrete precoding and finite-resolution RIS configurations. We transform the original sum rate maximization problem into an equivalent weighted sum MSE minimization problem and adopt an alternating optimization approach based on the block coordinate descent method to solve it.
  • We propose a solution based on the Schnorr Euchner SD (SESD) algorithm to solve the precoding optimization sub-problem. This solution provides the global optimal solution to the precoding optimization sub-problem while significantly reducing the complexity compared to an exhaustive search. We further employ the SESD algorithm to solve the RIS configuration optimization problem when RIS reflection coefficients belong to a discrete set. In contrast to the existing literature, where only sub-optimal solutions for discrete RIS configuration are offered, the SESD algorithm finds the optimal discrete RIS configurations. This is of paramount importance in a MU-MIMO RIS-assisted system, where sub-optimal configuration of the RIS limits the interference suppression capability of the system and severely degrades the performance.
@@
  • We evaluate all the proposed algorithms via numerical simulations under different system setups and show the superiority of the presented SESD-based designs over the commonly used benchmarks. We also assess the effectiveness of the proposed low-complexity heuristic SESD algorithm for RIS configuration optimization and show that it performs close to the optimal SESD design.
@@ III Iterative Weighted Sum MSE Minimization
in the sense that the optimal precoding $\mathbf{W}$ and optimal RIS configuration vector $\boldsymbol{\theta}$ are the same for both problems. In problem (9), $\mathbf{c}=[c_{1},\ldots,c_{K}]^{\mbox{\tiny$\mathrm{T}$}}$ and $\boldsymbol{\beta}=[\beta_{1},\ldots,\beta_{K}]^{\mbox{\tiny$\mathrm{T}$}}$.
@@ III-A Optimizing Precoding Vectors
$\mathbf{w}_{k}^{\mathrm{inf}}=\mathbf{V}^{-1}\mathbf{v}_{k}^{*}.$ (18)
@@
25:            $\tilde{\mathbf{w}}_{k}\leftarrow\check{\mathbf{w}}_{k}$
@@
$\xi_{m}=\tilde{\mathbf{d}}_{w,k}[m]-\tilde{\mathbf{R}}_{w}[m,m+1:2M]\check{\mathbf{w}}_{k}[m+1:2M].$ (24)
@@
is chosen as the new search radius. Then, fixing $\check{\mathbf{w}}_{k}[3],\ldots,\check{\mathbf{w}}_{k}[2M]$, the algorithm updates $\check{\mathbf{w}}_{k}[2]$ by selecting another label, i.e., $\mathcal{L}(\mathbf{p}_{2}[2])$ and checks if $\mathbf{r}[2]=\sum_{i=2}^{2M}\left|\tilde{\mathbf{d}}_{w,k}[i]-\tilde{\mathbf{R}}_{k}[i,i:2M]\tilde{\mathbf{w}}_{k}[i:2M]\right|^{2}$ with the newly updated $\check{\mathbf{w}}_{k}[2]$ exceeds the search radius $r_{\mathrm{opt}}$ or not. The reason why the algorithm does not update $\check{\mathbf{w}}_{k}[1]$ and directly goes for updating $\check{\mathbf{w}}_{k}[2]$ is that for the combination of entries $\check{\mathbf{w}}_{k}[2],\ldots,\check{\mathbf{w}}_{k}[2M]$, the value $\check{\mathbf{w}}_{k}[1]$ obtained by solving (23) for $m=1$ is the optimal choice among all the quantization labels and results in the smallest possible radius; thus, there is no need to explore other labels in the set at this level. If $\mathbf{r}[2]<r_{\mathrm{opt}}$, the algorithm proceeds to update $\check{\mathbf{w}}_{k}[1]$ by solving (23) again. Otherwise, it moves to an upper level and updates $\check{\mathbf{w}}_{k}[3]$ having $\check{\mathbf{w}}_{k}[4],\ldots,\check{\mathbf{w}}_{k}[2M]$ fixed. This procedure is repeated and after each successful update of $\check{\mathbf{w}}_{k}[m]$ (i.e., the update that results in $\mathbf{r}[m]<r_{\mathrm{opt}}$), the algorithm moves to a lower level and finds $\check{\mathbf{w}}_{k}[m-1]$ from (23). When updating $\check{\mathbf{w}}_{k}[m]$ is unsuccessful, the value of $m$ is increased by 1 and $\check{\mathbf{w}}_{k}[m+1]$ is updated by choosing another label from $\mathcal{L}$ based on the sorting indices at this level. Whenever a new combination of entries $\check{\mathbf{w}}_{k}[1],\ldots,\check{\mathbf{w}}_{k}[2M]$ is found that satisfies $\mathbf{r}[1]=\|\tilde{\mathbf{d}}_{w,k}-\tilde{\mathbf{R}}_{w}\check{\mathbf{w}}_{k}\|^{2}<r_{\mathrm{opt}}$, the search radius is updated as $r_{\mathrm{opt}}=\mathbf{r}[1]$. It is important to note the role of sorting indices in updating the precoding entries. Specifically, assuming that $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i])$, the next time the algorithm moves from level $m-1$ to level $m$, $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$ is selected as the new precoding entry at this level. If $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$ fails to fall in the optimal radius $r_{\mathrm{opt}}$, there is no need for further update of $\check{\mathbf{w}}_{k}[m]$ for the given values of $\check{\mathbf{w}}_{k}[m+1],\ldots,\check{\mathbf{w}}_{k}[2M]$. In other words, if $\mathbf{r}[m]>r_{\mathrm{opt}}$ for $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i+1])$, we will surely have $\mathbf{r}[m]>r_{\mathrm{opt}}$ for $\check{\mathbf{w}}_{k}[m]=\mathcal{L}(\mathbf{p}_{m}[i^{\prime}]),\,i^{\prime}>i+1$.
@@ III-B Optimizing the RIS Configuration
$\bar{\theta}_{\hat{n}}=e^{j\left(-\arg\left(\sum_{n\neq\hat{n}}\theta_{n}^{*}[\mathbf{A}]_{n,\hat{n}}-[\mathbf{t}]^{*}_{\hat{n}}\right)+\pi\right)}.$ (34)
@@ III-C Overall Algorithm
7:Given $\overline{\mathbf{W}}^{(0)}$, $\overline{\boldsymbol{\theta}}^{(0)}$, and $\overline{\boldsymbol{\beta}}^{0}$, find $e_{k}^{(0)}$ from (III) and compute $f^{(0)}=\sum_{k=1}^{K}\overline{c}_{k}^{(0)}e_{k}^{(0)}-\log_{2}(\overline{c}_{k}^{(0)})$
@@
10:   Using Algorithm 1 and bisection method, find the optimal precoding vectors $\{\overline{\mathbf{w}}_{k}^{(l)}\}$
@@
15:   Compute $f^{(l)}=\sum_{k=1}^{K}\overline{c}_{k}^{(l)}e_{k}^{(l)}-\log_{2}(\overline{c_{k}}^{(l)})$
@@
IV RIS with Finite Phase Shift Resolution
In practice, due to hardware design limitations, it may not be possible for RIS elements to take any phase value. Therefore, a practical RIS is expected to have a finite phase shift resolution, meaning that only a limited number of phase shifts are available for RIS elements to choose from. In such a case, the discrete set of reflection coefficients is given by